# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v4

@@ Abstract.
To tackle this obstacle we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a single optimization problem. By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; March xx–xx, 2025; California, USAisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code.
@@
Despite their advantages, existing HLS workflows face several significant challenges. Previous frameworks (Ye et al., 2022, 2024; Lai et al., 2019; Zhang et al., 2024) have integrated pragma insertion and code transformation, but these efforts are somewhat limited, focusing primarily on loop permutations based on loop properties. Additionally, code transformations and pragma insertion are often treated as separate processes. Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, while pragmas for optimizations like loop unrolling, pipelining, and array partitioning are applied later. This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa. Moreover, the immense design space—encompassing millions of potential transformation and pragma combinations—makes exhaustive exploration unfeasible. Ensuring the correctness of these transformations while adhering to resource constraints further complicates the process.
@@
Our primary research objective is to develop a system capable of autonomously conducting code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization framework. Even though several exploration methods could be used, we chose to employ a Nonlinear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
@@
  • A novel template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels. Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within seconds or minutes. Additionally, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP, all without the necessity of DSE. Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
The paper is organized into the following sections. Section 2 lays out the motivation behind our approach and the proposed solution. Following that, Section 3 presents the range of code transformation and pragma insertion we consider. In Section 4, we introduce a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Section 5 elaborates on our post-optimizations. Finally, sections 6, 7, 8 and 9 are dedicated to evaluating our method, presenting related work, addressing limitations and drawing conclusions.
@@ 2.1. Space to Explore
The various components of the exploration space—pragma insertion, code transformations, and data caching—are inherently interdependent. For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization. Similarly, the scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops. A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1). The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
Consequently, selecting one element impacts the others, often requiring backtracking when the spaces are treated separately. This interconnection adds complexity to the exploration process and necessitates ongoing verification to ensure the legality of transformations.
@@ 74 C[i][j]+=alpha*A[i][k2]*B[k2][j];}// S1
Listing 1: Original Code
@@ 32 C[i][j]+=alpha*A[i][k]*B[k][j];}//S1
Listing 2: Result found using the NLP with the constraints:
DSP = 6,840 On-chip mem = 7.2 MB
2.2. Limitation of the Current DSE
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024). Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning. Additionally, Pluto minimizes memory transfer from off-chip to on-chip (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat). Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging. Exploring every possible transformation is also impractical for large design spaces. The key differences between the frameworks are summarized in Table 1.
@@ 2.2. Limitation of the Current DSE
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a). A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6. AutoDSE treats the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD/Xilinx source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones. A more detailed description of the Merlin compiler can be found in (Pouget et al., 2024b). In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled. The targeted FPGA is the AMD/Xilinx Alveo U200. We evaluate two scenarios: the gemm kernel from Polybench (Polybench, [n. d.]), as shown in Listing 1, and a convolutional neural network (CNN) layer where the problem size and original loop order of CNN are I,J=256 (Output and Input Channels), H,W=224 (Height, Weight) and P,Q=5 (Filter Height and Weight). The gemm kernel’s original loop order, as illustrated in Listing 1, has a non-reduction loop at the center of the second statement (L3). This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop (L2). In this scenario, if the reduction loop is pipelined and partially unrolled by a factor of $uf$, performance will degrade. While partial unrolling allows for parallel execution of the multiplications, it increases the pipeline depth due to the loop’s dependencies, as the $uf$ additions must still be performed sequentially. Consequently, when the reduction loop is pipelined, the initiation interval (II) will significantly increase, becoming a multiple of both $uf$ and the latency of the reduction operation. This occurs because the next iteration of the loop can only begin after the $uf$ additions are completed and the output is fully written. Their inability to further increase parallelism due to loop order limits the best designs to a throughput of 20 GFLOPs per second (GF/s). In such instances, a transformation becomes imperative to augment parallelization. A single CNN layer with 6 loops results in an enormous design space that cannot be explored exhaustively. Loop permutations alone represent 6! = 720 possibilities, and tiling is required to manage array sizes exceeding on-chip memory capacity. While frameworks like AutoDSE and NLP-DSE achieve satisfactory throughput—42.15 GF/s and 31.80 GF/s, respectively—they are unable to explore the full design space, which includes both code transformations and pragma insertion, within a reasonable timeframe. For example, running AutoDSE for all possible loop permutations would take around 500 days, and for NLP-DSE, it would take 106 days. Other frameworks, which apply code transformations based solely on loop properties, such as prioritizing reduction loops in the outermost positions, explore a much more restricted design space, limiting their potential for further optimization. This underscores the challenge of navigating such a vast design space and highlights gaps in current approaches.
@@ 2.3. Overview of Sisyphus
To overcome these limitations, we introduce Sisyphus, a solution tailored for affine programs, where loop bounds and conditionals are affine functions of surrounding loop iterators and parameters (commonly referred to as Static Control Parts in the Polyhedral Model) (Girbal et al., 2006; Feautrier, 1988). Our objective is to define a search space that allows for the selection of various parameters, such as pragma insertion, loop ordering and tiling, within a single optimization problem. This approach avoids the need for backtracking or handling disjoint optimization problems. Our framework efficiently and rapidly explores this large design space, overcoming the limitations of other frameworks that either restrict the search space too much or explore a broad space with a DSE that only partially covers it. Most importantly, it guarantees that only valid code transformations are considered. To accomplish this, we propose a two-step process: first, the creation of a space tailored to each kernel, followed by the exploration of that space. The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations. To achieve this, we strip-mined each original loop twice, ensuring that one of these loops is included at each level (assuming the permutation is valid). As a result, each original loop can be utilized across all three levels. For instance, in Listing 1, the loop $k$ (L2) is used in level 0 alongside loop L5 and in level 2 with loop L11 in Listing LABEL:lst:small_mem. Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources. As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions. For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (line 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints. Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place. Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem. This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs. It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS, as done in AutoDSE, would be impractical due to the sheer size of the search space. In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods. For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop. In this configuration, the NLP identifies the setup shown in Listing 2. Both pipelined loops achieve an initiation interval (II) of 1, resulting in a throughput of 210 GFLOPs per seconds (GF/s). The reduction loop (L11) is partially unrolled, allowing the multiplications to be executed in parallel while performing four additions sequentially. This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential. For the CNN, the NLP selects all parameters within the design space, including loop order after exploring 6! possibilities, data transfer points, on-chip array size, and the level of parallelism. This comprehensive analysis takes just 13 minutes and yields a throughput of 341 GF/s. Thanks to a well-defined space that only considers legal transformations and exploration guided by a cost model formulated as an NLP, our framework is able to select the theoretical optimum in the entire space in a matter of seconds or minutes and generate a design.

### v5

# A Unified Framework for Automated Code Transformation and Pragma Insertion

@@ Abstract.
To tackle this obstacle, we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a common optimization framework. By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
@@
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; February 27–March 1, 2025; Monterey, CA, USAbooktitle: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays (FPGA ’25), February 27–March 1, 2025, Monterey, CA, USAdoi: 10.1145/3706628.3708873isbn: 979-8-4007-1396-5/25/02ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
@@ 1. Introduction
High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024; Cong et al., 2022). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024, 2025; Sohrabizadeh et al., 2022b, 2023, a; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code.
@@
Despite their advantages, existing HLS workflows face several significant challenges. Previous frameworks (Ye et al., 2022, 2024; Lai et al., 2019; Zhang et al., 2024) have integrated pragma insertion and code transformation, but these efforts are somewhat limited, focusing primarily on loop permutations based on loop properties. Additionally, code transformations and pragma insertion are often treated as separate processes. Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, followed by pragma insertion for optimization, like loop unrolling, pipelining and array partitioning. This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa. Moreover, the immense design space—encompassing millions of potential transformation and pragma combinations—makes exhaustive exploration unfeasible. Ensuring the correctness of these transformations while adhering to resource constraints further complicates the process.
@@
Our primary research objective is to develop a system capable of autonomously conducting code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization framework. Even though several exploration methods could be used, we chose to employ a Nonlinear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2025, 2024). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
@@
  • A novel optimization template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels. Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within minutes. Plus, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP – all without requiring DSE to perform multiple HLS compiler syntheses. Furthermore, Sisyphus identifies the correct transformation, even when no transformation is needed.
The paper is organized into the following sections. Section 2 explains our approach and solution. Following that, Section 3 covers code transformation and pragma insertion. Section 4 introduces a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Section 5 elaborates on our post-optimizations. Finally, sections 6, 7, 8 and 9 are dedicated to evaluating our method, presenting related work, addressing limitations and drawing conclusions.
@@ 2.1. Design Space to Explore
The various components of the exploration space—pragma insertion, code transformations, and data caching—are inherently interdependent. For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization. Similarly, loop scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops. A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1). The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
Selecting one optimization affects others, often requiring backtracking when spaces are handled separately. This interdependence complicates exploration and demands constant verification of transformation legality.
@@ 74 C[i][j]+=alpha*A[i][k2]*B[k2][j];}// S1
Listing 1: Original gemm Code
@@ 32 C[i][j]+=alpha*A[i][k]*B[k][j];}//S1
Listing 2: Result found for gemm kernel using the NLP with the constraints: DSP = 6,840 On-chip mem = 7.2 MB
2.2. Limitation of the Current DSE Methods
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024). Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning. Additionally, Pluto minimizes memory transfer from off-chip to on-chip communication (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat). Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2022b, 2023, a; Pouget et al., 2025) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging. Exploring every possible transformation is also impractical for large design spaces. The key differences between the frameworks are summarized in Table 1.
@@ 2.2. Limitation of the Current DSE Methods
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2022b) and NLP-DSE (Pouget et al., 2025, 2024). A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6. AutoDSE treats the source-to-source compiler Merlin (Xilinx, 2024) and the associated HLS tools as a black box, adjusting pragmas based on the identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones. A more detailed description of this compiler can be found in (Pouget et al., 2025). In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled. The targeted FPGA is the AMD/Xilinx Alveo U200. We evaluate two scenarios: the gemm kernel from Polybench (PolyBench, [n. d.]), as shown in Listing 1, and a convolutional neural network (CNN) layer where the problem size and original loop order of CNN are I,J=256 (Output and Input Channels), H,W=224 (Height, Weight) and P,Q=5 (Filter Height and Weight). The gemm kernel’s original loop order, as illustrated in Listing 1, has a non-reduction loop at the center of the second statement (L3). This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop (L2). In this scenario, if the reduction loop is pipelined and partially unrolled by a factor of $uf$, performance will degrade. While partial unrolling allows for parallel execution of the multiplications, it increases the pipeline depth due to the loop’s dependencies, as the $uf$ additions must still be performed sequentially. Consequently, when the reduction loop is pipelined, the initiation interval (II) will significantly increase, becoming a multiple of both $uf$ and the latency of the reduction operation. This occurs because the next iteration of the loop can only begin after the $uf$ additions are completed and the output is fully written. Their inability to further increase parallelism due to loop order limits the best designs to a throughput of 20 GFLOPs per second (GF/s). In such instances, a transformation becomes imperative to augment parallelization. A single CNN layer with 6 loops results in an enormous design space that cannot be explored exhaustively. Loop permutations alone represent 6! = 720 possibilities, and tiling is required to manage array sizes exceeding on-chip memory capacity. While frameworks like AutoDSE and NLP-DSE achieve satisfactory throughput—42.15 GF/s and 31.80 GF/s, respectively—they are unable to explore the full design space, which includes both code transformations and pragma insertion, within a reasonable timeframe. For example, running AutoDSE for all possible loop permutations would take around 500 days, and for NLP-DSE, it would take 106 days. Other frameworks, which apply code transformations based solely on loop properties, such as prioritizing reduction loops in the outermost positions, explore a much more restricted design space, limiting their potential for further optimization. This underscores the challenge of navigating such a vast design space and highlights gaps in current approaches.
@@ 2.3. Overview of Sisyphus
We introduce Sisyphus, tailored for affine programs with loop bounds and conditionals as affine functions (Girbal et al., 2006; Feautrier, 1988). It defines a unified search space for parameters like pragma insertion, loop ordering, and tiling, avoiding backtracking or disjoint optimizations. Sisyphus efficiently explores this design space, overcoming limitations of frameworks that either overly restrict or partially explore it. Crucially, it ensures only valid transformations. The process involves creating a kernel-specific space and exploring it. The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations. To achieve this, we strip-mined each original loop twice, ensuring that one of these loops is included at each level (assuming the permutation is valid). As a result, each original loop can be utilized across all three levels. For instance, in Listing 1, the loop $k$ (L2) is used in level 0 (L5) and in level 2 (L11) in Listing LABEL:lst:small_mem. Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources. As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions. For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (lines 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints. Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place. Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem as in NLP-DSE (Pouget et al., 2025). This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs. It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS (Sohrabizadeh et al., 2022b) would be impractical due to the sheer size of the search space. In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods. Sisyphus outperforms NLP-DSE (Pouget et al., 2025, 2024) by exploring a larger optimization space, including tiling and loop permutations, unlocking performance unattainable by NLP-DSE, which requires separate runs for each transformation. Automating the entire process, Sisyphus reduces exploration time and delivers superior results. It generates Vitis HLS pragmas with broader parameter support, like array partitioning, enhancing accuracy and design space coverage. Its high model precision enables the generation of at most one or two optimized designs, avoiding NLP-DSE’s extensive iterations. For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop. In this configuration, the NLP identifies the setup shown in Listing 2. Both pipelined loops achieve an initiation interval (II) of 1, resulting in a throughput of 210 GFLOPs per seconds (GF/s). The reduction loop (L11) is partially unrolled, allowing the multiplications to be executed in parallel while performing four additions sequentially. This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential. For the CNN, the NLP selects all parameters in the design space, including loop order (exploring 6! possibilities), on-chip array size, parallelism levels, etc. This analysis completes in 13 minutes, achieving a throughput of 341 GF/s. Thanks to a well-defined space that considers only legal transformations and exploration guided by a cost model formulated as an NLP, our framework can identify the theoretical optimum across the entire space and generate a design within seconds or minutes.